OpenStudy (anonymous):

Part 1 [3 points]: Find the solution to the equation 64^(3 + x) = 4^2x. Part 2 [4 points]: Use complete sentences to explain the method used to solve this equation.

OpenStudy (nikvist):

\[64^{3+x}=4^{2x}\]\[(4^3)^{3+x}=4^{2x}\]\[4^{3(3+x)}=4^{2x}\]\[3(3+x)=2x\]\[9+3x=2x\]\[x=-9\]